OpenStudy (anonymous):

Write both solutions separated by a comma: 2x^2+7x+5=0

OpenStudy (bahrom7893):

d = 49-40=9 x1 = (-7-3)/4=-10/4 x2 = (-7+3)/4=-1

OpenStudy (id21):

\[-10/4, -1\]

OpenStudy (bahrom7893):

-10/4, -1